OpenStudy (moazzam07):

Winnie used a probability simulator to roll a 6-sided number cube and flip a coin 100 times. The results are shown in the tables below:

OpenStudy (anonymous):

tables?

OpenStudy (moazzam07):

Using Winnie's simulation, what is the probability of rolling a 4 on the number cube and the coin landing heads up? 16 over 100 72 over 100 1232 over 10000 2432 over 10000

OpenStudy (anonymous):

what is the probability of rolling a 4 according to the table?

OpenStudy (anonymous):

28/100

OpenStudy (moazzam07):

ok

OpenStudy (anonymous):

the probability of landing heads up is 44/100

OpenStudy (anonymous):

so do 28 over 100 times 44 over 100

OpenStudy (moazzam07):

let me do it

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

I'll tell you if it is right

OpenStudy (moazzam07):

your too fast

OpenStudy (anonymous):

do you know how to multiply fractions?

OpenStudy (moazzam07):

77/627

OpenStudy (moazzam07):

625

OpenStudy (anonymous):



OpenStudy (anonymous):

44*28 is 1,232

OpenStudy (anonymous):

and 100*100 is 10,000

OpenStudy (anonymous):

did you use a fraction calc?

OpenStudy (anonymous):

just wondering...

OpenStudy (moazzam07):

yes :(

OpenStudy (anonymous):

I'll show you how to multiply fractions

OpenStudy (moazzam07):

I know how but i dont have paper on me

OpenStudy (moazzam07):

Give me a problem and ill answer it

OpenStudy (anonymous):

\[\frac{ 28 }{ 100 } \times \frac{ 44 }{ 100 } = \frac{ 1,232 }{ 10,000 }\]

OpenStudy (anonymous):

since 28*44 is 1232 and 100*100 is 10,000

OpenStudy (anonymous):

okay

OpenStudy (moazzam07):

ok

OpenStudy (anonymous):

do

OpenStudy (moazzam07):

do what?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }\times \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

do 3*1 across and 4*2 on the bottom

OpenStudy (moazzam07):

1 and 1/4

OpenStudy (moazzam07):

XD this is Sad

OpenStudy (anonymous):

no

OpenStudy (anonymous):

don't worry

OpenStudy (moazzam07):

3/8?

OpenStudy (anonymous):